# CreatiLayout: Siamese Multimodal Diffusion Transformer for Creative Layout-to-Image Generation

### v2

@@ Abstract
Diffusion models have been recognized for their ability to generate images that are not only visually appealing but also of high artistic quality. As a result, Layout-to-Image (L2I) generation has been proposed to leverage region-specific positions and descriptions to enable more precise and controllable generation. However, previous methods primarily focus on UNet-based models (e.g., SD1.5 and SDXL), and limited effort has explored Multimodal Diffusion Transformers (MM-DiTs), which have demonstrated powerful image generation capabilities. Enabling MM-DiT for layout-to-image generation seems straightforward but is challenging due to the complexity of how layout is introduced, integrated, and balanced among multiple modalities. To this end, we explore various network variants to efficiently incorporate layout guidance into MM-DiT, and ultimately present SiamLayout. To Inherit the advantages of MM-DiT, we use a separate set of network weights to process the layout, treating it as equally important as the image and text modalities. Meanwhile, to alleviate the competition among modalities, we decouple the image-layout interaction into a siamese branch alongside the image-text one and fuse them in the later stage. Moreover, we contribute a large-scale layout dataset, named LayoutSAM, which includes 2.7 million image-text pairs and 10.7 million entities. Each entity is annotated with a bounding box and a detailed description. We further construct the LayoutSAM-Eval benchmark as a comprehensive tool for evaluating the L2I generation quality. Finally, we introduce the Layout Designer, which taps into the potential of large language models in layout planning, transforming them into experts in layout generation and optimization.
Figure 1: We present a novel approach to empower MM-DiTs for layout-to-image generation, achieving high-quality and fine-grained controllable generation, as evidenced by the precise rendering of complex attributes (e.g. color, texture, shape, and text).
@@ 1 Introduction
Through comprehensive evaluations on LayoutSAM-Eval and COCO benchmarks, SiamLayout outperforms other variants and previous SOTA models by a clear margin, especially in generating entities with complex attributes, as illustrated in Fig. 1. For layout planning, LayoutDesigner shows more comprehensive and specialized capabilities compared to baseline LLMs.
@@ 2 Related Work
Text-to-image generation [55, 50, 3, 56, 6, 39, 48, 2, 14] has emerged as a promising application due to its impressive capabilities. Recently, studies such as SD3 [11], SD3.5 [61], FLUX.1 [33], and Playground-v3 [42] have advanced the multimodal Diffusion Transformer architecture (MM-DiT), elevating text-to-image generation to the next level. MM-DiT significantly enhances text understanding by treating text as an independent modality, equally important as the image, and replacing traditional cross-attention with MM-Attention for modal interaction.
@@
To achieve more precise and controllable generation, layout-to-image generation [38, 64, 74, 13, 9, 15, 59, 49, 35, 67, 28, 70, 73, 69, 22, 44, 7] has been proposed to generate images based on layout guidance, which includes several entities. Each entity comprises a spatial location and a region description. However, previous methods primarily focused on UNet-based architectures, and enabling MM-DiT for layout-to-image generation is challenging due to the complexity of introducing, integrating, and balancing the layout among multiple modalities. In this paper, we focus on exploring winner solutions for incorporating layout into MM-DiT, thereby fully unleashing its capabilities and achieving precisely controllable generation.
@@ 3.2 Layout-to-Image Generation
[…]
@@
  • Layout Adapter. Based on the fundamental idea of previous L2I methods [38, 64, 74, 72] introducing layout conditions in UNet-based architectures, we design extra image-layout cross-attention to incorporate layout into MM-DiT, defined as the Layout Adapter. Formally, it is represented as:
@@
    where $\mathbf{K}^{l}=\bm{h}^{l}\mathbf{W}_{k}^{l}$ and $\mathbf{V}^{l}=\bm{h}^{l}\mathbf{W}_{v}^{l}$ are the key and values matrices from the layout tokens. $\mathbf{W}_{k}^{l}$ and $\mathbf{W}_{v}^{l}$ are the corresponding weight matrices. $\mathbf{Q}^{z}$ is the identical query as in Eq. 7. In this way, guidance from the layout is introduced into $\bm{h}_{\mathrm{adapter}}^{z}$ with fewer parameters. However, due to layout conditions not playing an equal role to the global caption condition, it may diminish the impact of layout guidance on the generation process and lead to suboptimal results.
@@
    where ${\mathbf{Q}^{z}}^{\prime}$,${\mathbf{K}^{z}}^{\prime}$ and ${\mathbf{V}^{z}}^{\prime}$ are the new query, key, and value matrices obtained from the image token, with weight matrices different from those in Eq. 7. The final image tokens are the fusion of text-guided and layout-guided image tokens:$\bm{h}_{\mathrm{siamese}}^{z}={\bm{h}^{z}}^{\prime}+{\bm{h}^{z}}^{\prime\prime}$. Since the guidance from the layout and global caption has been decoupled—being independent in space and parallel in time—the issue of modality competition is significantly mitigated. As shown in Fig. 3 (b), in the attention map of image-layout MM-Attention, the similarity between layout and image increases continually during training as layout takes a dominant role here.
In this paper, SiamLayout is chosen as the network that incorporates layout into MM-DiTs, and our primary experiments are conducted on it. In addition, we develop a LoRA-based variant of SiamLayout, which achieves comparable layout control accuracy in a more lightweight manner. Further details can be found in the supplementary material.
Figure 3: Attention maps of $\mathbf{M}^{3}$-Attention and SiamLayout. (a) The layout’s influence on image generation is much lower compared to the text due to lower similarity. (b) SiamLayout decouples (a) into two independent MM-Attentions for image-text and image-layout, enabling equal guidance from both layout and text.
@@
Here, we employ two strategies to accelerate the convergence of the model: \@slowromancapi@) Biased sampling of time steps: Since layout pertains to the structural content of images, which is primarily generated during the larger time steps, we sample time steps with a 70% probability from a normal distribution $\mathcal{N}(0.7*T,T)$ and with a 30% probability from $\mathcal{N}(0,T)$. \@slowromancapii@) Region-aware loss: We enhance the model’s focus on areas specified by the layout by assigning greater weight to the region loss $\mathcal{L}_{region}$ associated with the regions localized by the bounding boxes in the latent space. The updated loss is:
@@ 3.3 Layout Dataset and Benchmark
  • Global-wise quality. This aspect scores based on visual quality and global caption following, across multiple metrics including recently proposed scoring models like IR score [68] and Pick score [32], as well as traditional metrics such as CLIP [52], FID [20] and IS [57] scores. For more details on the proposed dataset and benchmark, please refer to the supplementary materials.
@@ 3.4 Layout Designer
  • Generation for coarse-grained inputs. For user inputs with only a global caption, $\mathrm{M}_{l}$ designs from scratch based on the caption to determine which entities compose the layout and generates appropriate bounding boxes for these entities; for cases where the description and centroid coordinates ($x_{0}$,$y_{0}$) of each entity are provided, $\mathrm{M}_{l}$ designs harmonious bounding boxes based on this information.
@@ 4.2 Evaluation on Layout-to-Image Generation
Figure 5: Qualitative results on the LayoutSAM-Eval. SiamLayout outperforms previous methods by a clear margin, especially in generating complex attributes such as color, texture, text, and portrait details. Zoom in for details.
@@
AP $\uparrow$ $\mathrm{AP}^{50}$uparrow AR $\uparrow$ CLIP $\uparrow$ FID $\downarrow$ IS $\uparrow$
@@ 4.5 Ablation Study
We explore three network variants aimed at integrating layout guidance into MM-DiT. Tab. 6 illustrates the accuracy of layout adherence of these designs. Compared to the vanilla SD3, Layout Adapter enhances the model’s adherence to spatial locations and attributes by introducing cross-attention between the image and layout. However, as shown in Fig. 7, it falls short when dealing with complex and fine-grained color, quantity, and texture requirements. We attribute this to the fact that the layout is not considered an independent modality equally important as the global caption and image, which limits the potential of layout guidance. The initial intention behind designing $\mathrm{M}^{3}$-Attention is to make layout, image, and caption play an equal role. However, due to the competition among these modalities in the attention map of $\mathrm{M}^{3}$-Attention, layout modality consistently is at a disadvantage, which is reflected in both quantitative and qualitative results as lower layout responsiveness. SiamLayout decouples $\mathrm{M}^{3}$-Attention into two independent and parallel MM-Attention branches: image-text and image-layout. This design allows each branch to play a significant role without interfering with each other, jointly contributing to precise responses to spatial locations and complex attributes.
@@
In Fig. 8, we explore the impact of two strategies introduced during training on SiamLayout: biased time step sampling and region-aware loss. With the region-aware loss $\mathcal{L}_{region}$, the model focuses more on the areas localized by the layout, accelerating the model’s convergence. In addition, layout predominantly guides structural content, which is mainly generated at larger time steps. Thus, sampling larger time steps with a higher probability (i.e. biased time step sampling) also effectively speeds up the model’s convergence.
@@ 5 Conclusion
We presented SiamLayout, which treats layout as an independent modality and guides image generation through an image-layout branch that is siamese to the image-text branch. SiamLayout unleashes the power of MM-DiT to achieve high-quality, accurate layout-to-image generation, and significantly outperforms previous work in generating complex attributes such as color, shape, and text. Additionally, we introduced LayoutSAM, a large-scale layout dataset with 2.7M image-text pairs and 10.7M entities, along with LayoutSAM-Eval to assess generation quality. Finally, we proposed LayoutDesigner, which tames a large language model into a professional layout planner capable of handling user inputs of varying granularity.
@@
We introduce a large language model for layout planning, which brings extra computation costs. Integrating layout planning with layout-to-image generation into an end-to-end model is an important direction for future research. Additionally, the automatic annotation pipeline introduces noisy data mainly caused by the object detection model, so its impact on the performance of the layout-to-image model needs to be further studied.
@@ C Discussion on Extra Parameters and Computation Introduced by Layout Control
D User Study
Figure 11: The results of the user study.
To appropriately design the evaluation criteria for our user study, we solicited feedback from professional designers as well as a large number of general users. As shown in Fig. 11, the evaluation dimensions were eventually categorized into two groups: region-wise quality and global-wise quality. Specifically, region-wise quality assesses the accuracy of attributes such as spatial, color, shape, texture, text, and portrait details, while global quality evaluates the overall aesthetic quality, prompt-following capability, and reasonability of the generated images. We randomly sampled 100 examples to conduct the user study. The statistical results demonstrate that SiamLayout achieves clear superiority across multiple evaluation dimensions.
E More Details on Datasets and Benchmarks
Figure 12: An overview of the proposed annotation pipeline, which can automatically annotate the layout information for any given image.
E.1 LayoutSAM Dataset and Benchmark
@@
We design a mechanism to automatically annotate the layout for any given image, as shown in Fig. 12.
@@
E.2 Layout Planning Dataset and Benchmark
@@
F More Analysis on Modal Competition
@@
G More Qualitative Results

### v3

@@ Abstract
Corresponding author.
Diffusion models have been recognized for their ability to generate images that are not only visually appealing but also of high artistic quality. As a result, Layout-to-Image (L2I) generation has been proposed to leverage region-specific positions and descriptions to enable more precise and controllable generation. However, previous methods primarily focus on UNet-based models (e.g. SD1.5 and SDXL), and limited effort has explored Multimodal Diffusion Transformers (MM-DiTs), which have demonstrated powerful image generation capabilities. Enabling MM-DiT for layout-to-image generation seems straightforward but is challenging due to the complexity of how layout is introduced, integrated, and balanced among multiple modalities. To this end, we explore various network variants to efficiently incorporate layout guidance into MM-DiT, and ultimately present SiamLayout. To inherit the advantages of MM-DiT, we use a separate set of network weights to process the layout, treating it as equally important as the image and text modalities. Meanwhile, to alleviate the competition among modalities, we decouple the image-layout interaction into a siamese branch alongside the image-text one and fuse them in the later stage. Moreover, we contribute a large-scale layout dataset, named LayoutSAM, which includes 2.7 million image-text pairs and 10.7 million entities. Each entity is annotated with a bounding box and a detailed description. We further construct the LayoutSAM-Eval benchmark as a comprehensive tool for evaluating the L2I generation quality. Finally, we introduce the Layout Designer, which taps into the potential of large language models in layout planning, transforming them into experts in layout generation and optimization. These components form CreatiLayout—a systematic solution that integrates the layout model, dataset, and planner for creative layout-to-image generation. Code, model, and dataset:https://creatilayout.github.io/.
Figure 1: We present a novel approach to empower $M$-DiTs for layout-to-image generation, achieving high-quality and fine-grained controllable generation, as evidenced by the precise rendering of complex attributes (e.g. color, texture, shape, and text).
@@ 1 Introduction
These three components—SiamLayout, LayoutSAM, and LayoutDesigner—jointly form the systematic layout-to-image solution CreatiLayout, which integrates the layout model, dataset, and planner. Through comprehensive evaluations on LayoutSAM-Eval and COCO benchmarks, SiamLayout outperforms other variants and previous SOTA models by a clear margin, especially in generating entities with complex attributes, as illustrated in Fig. 1. For layout planning, LayoutDesigner shows more comprehensive and specialized capabilities compared to baseline LLMs.
@@ 2 Related Work
Text-to-image generation [57, 52, 3, 58, 6, 40, 50, 2, 14, 74, 66] has emerged as a promising application due to its impressive capabilities. Recently, studies such as SD3 [11], SD3.5 [63], FLUX.1 [33], and Playground-v3 [43] have advanced the multimodal Diffusion Transformer architecture (MM-DiT), elevating text-to-image generation to the next level. MM-DiT significantly enhances text understanding by treating text as an independent modality, equally important as the image, and replacing traditional cross-attention with MM-Attention for modal interaction.
@@
To achieve more precise and controllable generation, layout-to-image generation [39, 67, 81, 13, 9, 15, 61, 51, 35, 70, 28, 77, 80, 71, 73, 72, 76, 22, 45, 7, 48, 37] has been proposed to generate images based on layout guidance, which includes several entities. Each entity comprises a spatial location and a region description. However, previous methods primarily focused on UNet-based architectures, and enabling MM-DiT for L2I generation is challenging due to the complexity of introducing, integrating, and balancing the layout among multiple modalities. In this paper, we focus on exploring winner solutions for incorporating layout into MM-DiT to unlock its power and enable precise and controllable generation.
@@ 3.2 Layout-to-Image Generation
$\text{Layout Condition: }\bm{l}=[\bm{e}_{1},\cdots,\bm{e}_{N}],$ $\mathrm{with}$ (4)
$\text{Entity: }\bm{e}_{n}=(c_{n},b_{n}).$ (5)
@@
  • Layout Adapter. Based on the fundamental idea of previous L2I methods [39, 79] introducing layout conditions in UNet-based architectures, we design extra image-layout cross-attention to incorporate layout into MM-DiT, defined as the Layout Adapter:
@@
    where $\mathbf{K}^{l}=\bm{h}^{l}\mathbf{W}_{k}^{l}$ and $\mathbf{V}^{l}=\bm{h}^{l}\mathbf{W}_{v}^{l}$ are the key and values matrices from the layout tokens. $\mathbf{W}_{k}^{l}$ and $\mathbf{W}_{v}^{l}$ are the corresponding weight matrices. $\mathbf{Q}^{z}$ is the identical query as in Eq. 7. Guidance from the layout is introduced into $\bm{h}_{\mathrm{adapter}}^{z}$ with fewer parameters. However, due to layout conditions not playing an equal role to the caption condition, it may diminish the impact of layout guidance on the generation process and lead to suboptimal results.
@@
[…]
In this paper, SiamLayout is chosen as the network that incorporates layout into MM-DiTs, and our primary experiments are conducted on it. In addition, we develop a LoRA-based variant of SiamLayout, which achieves comparable layout control accuracy in a more lightweight manner.
Figure 3: Attention maps of $\mathbf{M}^{3}$-Attention and SiamLayout. (a) The layout’s influence on image generation is much lower compared to the text due to lower similarity. (b) SiamLayout decouples (a) into two independent $M$-Attentions for image-text and image-layout, enabling equal guidance from both layout and text.
@@
Here, we employ two strategies to accelerate the convergence of the model: \@slowromancapi@) Biased sampling of time steps: Since layout pertains to the structural content of images, which is primarily generated during the larger time steps, we sample time steps with a 70% probability from a normal distribution $\mathcal{N}(0.7*T,T)$ and with a 30% probability from $\mathcal{N}(0,T)$. \@slowromancapii@) Region-aware loss: We enhance the model’s focus on areas specified by the layout by assigning greater weight to the region loss $\mathcal{L}_{region}$ associated with the regions localized by the bounding boxes in the latent space. The updated loss:
@@ 3.3 Layout Dataset and Benchmark
  • Global-wise quality. This aspect scores based on visual quality and global caption following, across multiple metrics including recently proposed scoring models like IR score [75] and Pick score [32], as well as traditional metrics such as CLIP [54], FID [20] and IS [59] scores.
@@ 3.4 Layout Designer
  • Generation for coarse-grained inputs. For user inputs with only a global caption, $\mathrm{M}_{l}$ designs from scratch based on the caption to determine which entities compose the layout and generates appropriate bounding boxes for these entities; for cases where the description and centroid coordinates ($x_{0}$$y_{0}$) of each entity are provided, $\mathrm{M}_{l}$ designs harmonious bounding boxes based on this information.
@@ 4.2 Evaluation on Layout-to-Image Generation
Figure 6: Qualitative results on the LayoutSAM-Eval. SiamLayout outperforms previous methods by a clear margin, especially in generating complex attributes such as color, texture, text, and portrait.
@@
AP $\uparrow$ $\mathrm{AP}^{50}$ $\uparrow$ AR $\uparrow$ CLIP $\uparrow$ FID $\downarrow$ IS $\uparrow$
@@ 4.5 Ablation Study
We explore three network variants aimed at integrating layout guidance into MM-DiT. Tab. 6 illustrates the accuracy of layout adherence of these designs. Compared to the vanilla SD3, Layout Adapter enhances the model’s adherence to spatial locations and attributes by introducing cross-attention between the image and layout. However, as shown in Fig. 8, it falls short when dealing with complex color, quantity, and texture requirements. We attribute this to the fact that the layout is not considered an independent modality equally important as the global caption and image, which limits the potential of layout guidance. The initial intention behind designing $\mathrm{M}^{3}$-Attention is to make layout, image, and caption play an equal role. However, due to the competition among these modalities in the attention map of $\mathrm{M}^{3}$-Attention, layout modality consistently is at a disadvantage, which is reflected in both quantitative and qualitative results as lower layout responsiveness. SiamLayout decouples $\mathrm{M}^{3}$-Attention into two parallel MM-Attention branches: image-text and image-layout. This design allows each branch to play an independent role without interfering with each other, jointly contributing to precise responses to spatial locations and complex attributes.
@@
In Fig. 9, we explore the impact of two strategies introduced during training on SiamLayout: biased time step sampling and region-aware loss. With the region-aware loss $\mathcal{L}_{region}$, the model focuses more on the areas localized by the layout, accelerating the model’s convergence. In addition, since layout mainly guides structural content generated at larger time steps, sampling these steps with higher probability (i.e. biased time step sampling) also speeds up convergence.
@@ 5 Conclusion
We presented CreatiLayout, a layout-to-image (L2I) system that integrates three key components: a layout model (SiamLayout), a dataset (LayoutSAM), and a planner (LayoutDesigner). SiamLayout treats layout as an independent modality and guides image through an image-layout branch that is siamese to the image-text branch, unleashing the power of MM-DiT for high-quality and accurate L2I generation. It outperforms previous work in generating complex attributes such as color, shape, and text. LayoutSAM is a large-scale layout dataset with 2.7M image-text pairs and 10.7M entities, accompanied by LayoutSAM-Eval for comprehensive evaluation of quality. LayoutDesigner leverages a large language model as a professional layout planner, capable of handling diverse user inputs with varying granularity.
@@
Introducing a large language model for layout planning brings extra computation costs. Integrating layout planning with L2I generation into an end-to-end model is an important direction for future research. In addition, improving the accuracy of control in layouts with a large number of entities, or in scenarios where entities contain rich text, is also a promising direction.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (Grant 62032006 and Grant 62472098) and ByteDance (No.CT20230914000147).
@@ C Discussion on Extra Parameters and Computation Introduced by Layout Control
D More Details on Datasets and Benchmarks
D.1 LayoutSAM Dataset and Benchmark
@@
We design a mechanism to automatically annotate the layout for any given image.
@@
D.2 Layout Planning Dataset and Benchmark
@@
E More Analysis on Modal Competition
@@
F More Qualitative Results